# Rethinking Weight-Averaged Model-merging

### v1

@@ 1 Introduction
  • An examination of the patterns contained in the averaged model weights through the visualization of the patterns present in both linear layer as classifiers or inside deep learning models used on multiple datasets. Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
@@
  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of of model regularization.
@@ 2 Related Work
In recent years, model-merging [26, 17, 21, 4, 2, 19, 23, 35] has attracted growing interest. This approach has led to a range of methods focused on averaging or merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [33] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup”. Further, Yadav et al. proposed Ties-Merging [34] based on parameter averaging by considering conflicts between the model weights during training, allowing for smoother parameter alignment across models for merging. Some other approaches [1, 25] address the problem from a different perspective, focusing on finding weight merging operations via the similarity analyses of model features. Yu et al. [39] combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a more thorough analysis of weight-averaged model-merging than the study in [33], which serves as the foundation work for the area.
@@ 3 The Patterns Contained in Model Weights
In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching. Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with inner product which is a widely explored operation in neural networks. By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
@@ 3.1 Connections Between Inner Product and Template Matching
In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in the Fig. 6.1. A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image’s feature representation and each class’s template. The class whose template yields the highest activation is selected as the predicted class, indicating that the template matches the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different “important” features that contribute the most towards the objectives.
This inner product computation is essentially a similarity measurement. In this context, the averaging operation in the weight space represents a linear combination of weight matrices (shown as Fig. 2), which can be seen as Mixup [40] operation in the weight space. The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space. This also explains why recent research [5, 41, 10] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
This concept extends to model-merging, where multiple models or variations of a model are combined. In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model. The model after merging can work more effectively than individuals because they represent a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
@@ 4.2 Impact of Non-linear Activation Functions
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analysis above and the experimental study shown in Sec. 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights. However, the averaging on weights can show relatively good performance without introducing extra inference computational cost. Hence, this forms a trade-off between computation usage and model performance when deciding between model-merging or model ensemble.
@@ 5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
@@ 5.1 Analysis of Weight Magnitude After Averaging
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$. This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound equal to the maximum magnitude of the original weights, but in most cases it will be less than that maximum. Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
@@ 5.2 Variance Reduction in Model Weight Averaging
Mathematical Justification. We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The variances for model A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively. If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$. This shows that, when the variances of the two models are equal, the variance is reduced by a factor of two. More generally, since the average of the variances is always less than or equal to the maximum variance, we have:
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of decision boundary.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec. 6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximun magnitudes/variances of individual models before merging. This is because generally the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness. In this case, model-merging is similar to regularization. This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes may improve performance. However, for tasks requiring detailed pattern recognition, lower weight magnitudes may hinder the network’s ability to capture critical pattern distinctions in the data.
@@ 6.1 The Patterns Contained in Model Weights
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 clearly shows the patterns of the visual class learned, which are strongly correlated with the average of all images within each class. The template of each class acts as a merged collection of representative images of the respective class. Interestingly, if we merge the weights of first 5 classes in Fig. 1 with the last 5 classes, it gives us the results of Fig. 2. As we analyzed before, the weight-averaged model-merging acts as the Mixup augmentation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that after merging, each element within the combined template becomes noticeably more faded. This phenomenon can also explain the limitation of expressiveness of merged model as described in Sec. 4.
@@
Fig. 6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes. For example, “goldfish” shows a redish center in the averaged images and also shows a redish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visuliazation. However, they are not as clear as for Cifar100. This may due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear. This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [8] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection being similar to the Haar features described in Sec. 3.1.
@@
The visualization for every class in Cifar100 and Tiny ImageNet are put in the supplementary materials.
6.2 Averaging on Weights vs. Averaging on Features
6.2.1 Differences over Different Models
@@
As shown in the Fig. 3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst. We can also discover that greedy soups are generally better than uniform soups due to its mechanism to greedily select the best performing model to merge. It ensures that a good result can be reached by only incorporating the first model. Greedy ensemble on both logits and features also produce results that are comparable to greedy soups.
@@
In general, adding more models into the merging process, leads to better performance for all methods, including model-merging and ensemble. Interestingly, the ViT model, including 3 ViT based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups. This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs). CNNs is more adapted to the learning of local patterns within different scales of images, making it ideal for model-merging. On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model-merging. ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
@@
We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (as a function of the number of merged models and different model architectures, but excluding ViT models as their gaps are too large) in Fig. 8. We can see that the more models into the merging process, the larger the gaps between logits ensemble and uniform soups. The gaps are different according to different models, but generally there is down trend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm). This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities. Batch normalization can break this down trend. We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that the greedy soups can alleviate this marginal effect.
6.2.2 Differences over Different Datasets
@@
For the dataset-wise examination, we perform model merging and ensemble on 7 prevalent datasets with different data sizes. From Tab. 1, we can observe ensemble still generally performs better than merging. Different datasets have different performances and this is due to the characteristics of each dataset. Similarly to the model-wise comparison, with more models in the merging (2-7 models merging results in the supplementary materials), the better the performance for both merging and ensembling. The model ensembling is more stable than model merging in terms of the best and worst performance on different data. For example, on the PathMNIST [37, 38] dataset, uniform soup merging does not work really well compared to the relative performance on the Cifar10/Cifar100 or Tiny ImageNet datasets (that only achieves 40.85% accuracy, which is largely behind model ensemble methods). A similar situation happens on Celeba [15] dataset: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model. These results may be caused by the contradictions between two models’ weights.
@@ 6.2.3 Merging Models trained over different Tasks
Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods. Model 1 means VGG19 model trained on Cifar100 dataset; while Model 2 denotes the VGG19 models trained on each different dataset with different task. Ft ens is better than logit ens.
@@
We also explore model merging between different well-trained models on different tasks, as shown in Tab. 2. Generally, merging within Tiny ImageNet (where Model 2 was trained on) has the best performance of 61.02% accuracy, which is better than the ensemble on logits (58.40% accuracy). This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images. Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results does not perform better than model ensemble, as it is too small. On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement. This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data were learned. However, the performance on Celeba is not good. This may be caused by the Celeba dataset only containing human faces, which can be seen as a fine-grained classification task. Due to the uniformity of the data format within Celeba, it contradicts the diverse and rich general computer vision features learned by the model weights on Cifar100.

### v2

@@ 1 Introduction
  • An examination of the patterns contained in the model weights through the visualization of the patterns present in both linear layers as classifiers or inside deep learning models used on multiple datasets. Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
@@
  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of model regularization.
@@ 2 Related Work
In recent years, model-merging [26, 17, 21, 4, 2, 19, 23, 35] has attracted growing interest. This approach has led to a range of methods focused on averaging or merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [33] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup”. Further, Yadav et al. proposed Ties-Merging [34] based on parameter averaging by considering conflicts between the model weights during training, allowing for smoother parameter alignment across models for merging. Some other approaches [1, 25] address the problem from a different perspective, focusing on finding weight merging operations via the similarity analyses of model features. Yu et al. [39] combine models trained on domain-specific data to create a domain-agnostic model merging that generalizes well across diverse distributions.
Despite these advances, research into the underlying mechanisms that make model merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a more thorough analysis of weight-averaged model-merging than the study in [33], which serves as the foundation work for the area.
@@ 3 The Patterns Contained in Model Weights
In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching. Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks. By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
@@ 3.1 Connections Between Inner Product and Template Matching
In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in Fig. 6.1. A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image’s feature representation and each class’s template. The class whose template yields the highest activation is selected as the predicted class, indicating that the template matches the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different “important” features that contribute the most towards the objectives.
This inner product computation is essentially a similarity measurement. In this context, the averaging operation in the weight space represents a linear combination of weight matrices (shown as Fig. 2), which can be seen as Mixup [40] operation in the weight space. The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space. This also explains why recent research [5, 41, 10, 44] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
This concept extends to model merging, where multiple models or variations of a model are combined. In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model. The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
@@ 4.2 Impact of Non-linear Activation Functions
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analysis above and the experimental study shown in Sec. 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights. However, the averaging on weights can show relatively good performance without introducing extra inference computational costs. Hence, this forms a trade-off between computation usage and model performance when deciding between model merging or model ensemble.
@@ 5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes with VGG19 model on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
@@ 5.1 Analysis of Weight Magnitude After Averaging
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$. This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound is equal to the maximum magnitude of the original weights, but in most cases, it will be less than that maximum. Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
@@ 5.2 Variance Reduction in Model Weight Averaging
Mathematical Justification. We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The variances for models A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively. If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$. This shows that, when the variances of the two models are equal, the variance is reduced by a factor of two. More generally, since the average of the variances is always less than or equal to the maximum variance, we have:
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of the decision boundary.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec. 6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging. This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness. In this case, model merging is similar to regularization. This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes may improve performance. However, for tasks requiring detailed pattern recognition, lower weight magnitudes may hinder the network’s ability to capture critical pattern distinctions in the data.
@@ 6.1 The Patterns Contained in Model Weights
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 clearly shows the patterns of the visual class learned, which are strongly correlated with the average of all images within each class. The template of each class acts as a merged collection of representative images of the respective class. Interestingly, if we merge the weights of the first 5 classes in Fig. 1 with the last 5 classes, it gives us the results of Fig. 2. As we analyzed before, the weight-averaged model-merging acts as the Mixup augmentation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that after merging, each element within the combined template noticeably fades. This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec. 4.
@@
Fig. 6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes. For example, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visualization. However, they are not as clear as for Cifar100. This may be due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear. This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [8] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec. 3.1.
@@
The visualization for every class in Cifar100 and Tiny ImageNet is put in the supplementary materials.
6.2 Weights vs. Features Averaging
6.2.1 Observations over Different Models
@@
As shown in Fig. 3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst. We can also discover that greedy soups are generally better than uniform soups due to their mechanism to greedily select the best performing model to merge. It ensures that a good result can be reached by only incorporating the first model. The greedy ensemble on both logits and features also produces results that are comparable to greedy soups.
@@
In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensemble. Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups. This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs). CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for model merging. On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model merging. ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
@@
We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (as a function of the number of merged models and different model architectures, but excluding ViT models as their gaps are too large) in Fig. 8. We can see that the more models in the merging process, the larger the gaps between logits ensemble and uniform soups. The gaps are different according to different models, but generally, there is a downtrend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm). This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities. Batch normalization can break this downtrend. We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that greedy soups can alleviate this marginal effect.
6.2.2 Observations over Different Datasets
@@
For the dataset-wise examination, we perform model merging and ensemble on 7 prevalent datasets with different data sizes. From Tab. 1, we can observe ensemble still generally performs better than merging. Different datasets have different performances and this is due to the characteristics of each dataset. Similarly to the model-wise comparison, with more models in the merging (2-7 models merging results in the supplementary materials), the better the performance for both merging and ensembling. The model ensembling is more stable than model merging in terms of the best and worst performance on different data. For example, on the PathMNIST [37, 38] dataset, uniform soup merging does not work really well compared to the relative performance on the Cifar10/Cifar100 or Tiny ImageNet datasets (that only achieves 40.85% accuracy, which is largely behind model ensemble methods). A similar situation happens on Celeba [15] dataset: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model. These results may be caused by the contradictions between the two models’ weights.
@@ 6.2.3 Merging Models trained over different Tasks
Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods. Model 1 is the VGG19 model trained on the Cifar100 dataset, while Model 2 denotes the VGG19 model trained on each different dataset with different tasks. Feature ensemble is better than logits ensemble.
@@
We also explore model merging between different well-trained models on different tasks, as shown in Tab. 2. Generally, merging within Tiny ImageNet (where Model 2 was trained on) has the best performance of 61.02% accuracy, which is better than the ensemble on logits (58.40% accuracy). This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images. Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results do not perform better than the model ensemble, as it is too small. On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement. This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data was learned. However, the performance on Celeba is not good. This may be caused by the Celeba dataset only containing human faces, which can be seen as a fine-grained classification task. Due to the uniformity of the data format within Celeba, it contradicts the diverse and rich general computer vision features learned by the model weights on Cifar100.